# Onsager-Machlup Functional and Large Deviation Principle for Stochastic Hamiltonian Systems

### v1

@@ 6. Preservation of Invariant Tori in Nearly Integrable Stochastic Hamiltonian Systems
(6.6) $\sup_{D_{r_{j},s_{j}}(\mathscr{D}_{j})}|\text{Im}\phi^{j}|\leq\frac{\xi_{j}}{2}.$
Hence,
@@
Thus, thanks to (6.5), $\mathcal{H}^{\epsilon}_{j}\circ\phi^{j}=\mathcal{H}^{j}+\mathcal{P}^{j}$ satisfies the assumptions in (2.1) with $\varepsilon\sim\left\lVert\mathcal{P}^{j}\right\rVert_{r_{j},s_{j},\mathscr{D}_{j}}$, $r\sim r_{j}$, $s\sim s_{j}$, $\sigma\sim\sigma_{j}$, $C_{H}\sim 2C_{H}$ as
@@
Hence, in order to apply Lemma 6 to $\mathcal{H}^{\epsilon}_{j}\circ\phi^{j}=\mathcal{H}^{j}+\mathcal{P}^{j}$, we only need to check (2.3). Upon observation, we have
@@
Furthermore, based on Lemmas 2.15 and 2.16, we can obtain the convergence results for $G^{j}$, $P_{j}$, $\phi^{j}$, and $H_{j}$ as follows:
  • The sequence $G^{j}:=G_{j}\circ G_{j-1}\circ\cdots\circ G_{2}\circ G_{1}$ converges uniformly on $\mathscr{D}_{0}$ to a diffeomorphism $G_{*}:\mathscr{D}_{0}\to\mathscr{D}_{*}:=G_{*}(\mathscr{D}_{0})\subset\mathscr{D}$, and $G_{*}\in C_{W}^{1}(\mathscr{D}_{0})$.
  • $P_{j}$ converges uniformly to 0 on $\mathscr{D}_{*}\times\mathbb{T}_{s_{*}}^{d}$ in the $C_{W}^{2}$ topology.
@@
    with $\phi_{*}\in C_{W}^{\tilde{m}}(\mathscr{D}_{*}\times\mathbb{T}^{d})$ and $\phi_{*}(\cdot,\cdot)\in C^{\tilde{m}v}(\mathbb{T}^{d})$, for any given $y_{*}\in\mathscr{D}_{*}$.
@@
Based on the preceding proof, we can demonstrate that there exists a Cantor-like set $\mathcal{D}_{*}\subset\mathcal{D}$, an embedding $\phi_{*}=(v_{*},u_{*}):\mathcal{D}_{*}\times\mathbb{T}^{d}\rightarrow\mathcal{H}:=\phi_{*}(\mathcal{D}_{*}\times\mathbb{T}^{d})\subset\mathcal{D}\times\mathbb{T}^{d}$ of class $C_{W}^{\beta}(\mathcal{D}_{*}\times\mathbb{T}^{d})$, and a function $H_{*}\in C_{W}^{2}(\mathcal{D}_{*},\mathbb{R})$, such that $H^{\epsilon}\circ\phi_{*}(I_{*},\theta)=K_{*}(I_{*},\theta)$ for all $(I_{*},\theta)\in\mathcal{D}_{*}\times\mathbb{T}^{d}$. The map $\xi\mapsto\phi_{*}(I_{*},\theta)$ is of class $C_{\beta}^{v}(\mathbb{T}^{d})$ for any $I_{*}\in\mathcal{D}_{*}$ (with $v^{-1}<\beta<\beta_{0}$), and the map $G^{*}:=(\partial_{I_{*}}H_{*})^{-1}\circ\partial_{I}H:\mathcal{D}_{*}\rightarrow\mathcal{D}_{*}$ defines a lipeomorphism onto $\mathcal{D}_{*}$, satisfying $B_{\alpha/2}(\mathcal{D}_{*})\subseteq\mathcal{D}$. The set $\mathcal{K}$ is foliated by KAM tori of $H^{\epsilon}$, each being a graph of a $C^{v}(\mathbb{T}^{d})$-map.
@@
$\epsilon^{2}\ln\mathbb{P}(X(t)\in A)\approx-\inf_{\varphi\in A}J(\varphi),$
where $\varphi-x_{0}\in\mathbb{H}^{1}$, $A\in\mathbb{R}^{d}$ denotes an arbitrary measurable set and the rate function $J(\varphi)$ is given by:
$J(\varphi)$ $=\frac{1}{2}\left(\int_{0}^{1}\left\|\sigma_{\theta}^{-1}(t)\left(\dot{\varphi}_{\theta}-\frac{\partial H_{0}}{\partial\varphi_{I}}(\varphi_{\theta},\varphi_{I})\right)\right\|^{2}\,{\rm d}t\right.$
$\qquad\left.+\int_{0}^{1}\left\|\sigma_{I}^{-1}(t)\left(\dot{\varphi}_{I}+\frac{\partial H_{0}}{\partial\varphi_{\theta}}(\varphi_{\theta},\varphi_{I})\right)\right\|^{2}\,{\rm d}t\right),$
where $\epsilon_{2}$ represents the strength of the stochastic perturbation, $\sigma_{\theta}^{-1}(t)$ and $\sigma_{I}^{-1}(t)$ are the inverses of the diffusion matrices $\sigma_{\theta}(t)$ and $\sigma_{I}(t)$, respectively. ∎
@@
$\mathbb{P}(X_{\epsilon}(t)\in\mathcal{K}_{0})\approx\exp\left\{-C\frac{\epsilon_{1}^{2}}{\epsilon_{2}^{2}}\times\mathbb{P}(\left\|W^{\sigma}\right\|\leq 1)\right\}\times\mathbb{P}(\left\|W^{\sigma}\right\|\leq 1).$
@@
Based on the content of Theorem 6.1, we have:
$\quad\frac{\mathbb{P}(X_{\epsilon}(t)\in\mathcal{K}_{0})}{\mathbb{P}(\left\|W^{\sigma}\right\|\leq 1}\approx\exp\left\{-\frac{1}{\epsilon_{2}^{2}}\inf_{\varphi\in\mathcal{K}_{0}}I(\varphi)\right\}$
$=\exp\left\{-\frac{1}{2\epsilon_{2}^{2}}\inf_{\varphi\in\mathcal{K}_{0}}\left(\int_{0}^{1}\left\|\sigma_{q}^{-1}(t)\left(\epsilon_{1}\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\right)\right\|^{2}\,{\rm d}t\right.\right.$
$\qquad\qquad\qquad\qquad\left.\left.+\int_{0}^{1}\left\|\sigma_{p}^{-1}(t)\left(\epsilon_{1}\frac{\partial H}{\partial\varphi_{q}}(\varphi_{q},\varphi_{p})\right)\right\|^{2}\,{\rm d}t\right)\right\}$
$=\exp\left\{-\frac{\epsilon_{1}^{2}}{2\epsilon_{2}^{2}}\inf_{\varphi\in\mathcal{K}_{0}}\left(\int_{0}^{1}\left\|\sigma_{q}^{-1}(t)\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\right\|^{2}\,{\rm d}t\right.\right.$
$\qquad\qquad\qquad\qquad\left.\left.+\int_{0}^{1}\left\|\sigma_{p}^{-1}(t)\frac{\partial H}{\partial\varphi_{q}}(\varphi_{q},\varphi_{p})\right\|^{2}\,{\rm d}t\right)\right\}$
@@
where $C$ is a quantity that depends on $\sigma_{p}^{-1}$, $\sigma_{q}^{-1}$ and the partial derivatives of $H$. ∎
@@
(6.7) $\begin{cases}\mathrm{d}q_{1}(t)=p_{1}(t)\,\mathrm{d}t-\epsilon p_{2}(t)\,\mathrm{d}t+\epsilon(2+\sin(t))\,\mathrm{d}W_{1}(t),\\ \mathrm{d}p_{1}(t)=-2q_{1}(t)\,\mathrm{d}t+\epsilon\sin(0.6t)\,\mathrm{d}t+\epsilon(2+\cos(t))\,\mathrm{d}W_{2}(t),\\ \mathrm{d}q_{2}(t)=p_{2}(t)\,\mathrm{d}t-\epsilon p_{1}(t)\,\mathrm{d}t+\epsilon(1+2\sin(t))\,\mathrm{d}W_{3}(t),\\ \mathrm{d}p_{2}(t)=-q_{2}(t)\,\mathrm{d}t+\epsilon\cos(0.6t)\,\mathrm{d}t+\epsilon(1+2\cos(t))\,\mathrm{d}W_{4}(t).\end{cases}$
@@
(6.8) $\begin{cases}\mathrm{d}q_{1}(t)=p_{1}(t)\,\mathrm{d}t-\epsilon p_{2}(t)\,\mathrm{d}t,\\ \mathrm{d}p_{1}(t)=-2q_{1}(t)\,\mathrm{d}t+\epsilon\sin(0.6t)\,\mathrm{d}t,\\ \mathrm{d}q_{2}(t)=p_{2}(t)\,\mathrm{d}t-\epsilon p_{1}(t)\,\mathrm{d}t,\\ \mathrm{d}p_{2}(t)=-q_{2}(t)\,\mathrm{d}t+\epsilon\cos(0.6t)\,\mathrm{d}t.\end{cases}$
@@
The Hamiltonian of this system can be written as:
@@
From the theorems discussed earlier, we know that the most probable continuous path of the stochastic nearly integrable Hamiltonian system (6.7) is governed by the deterministic equations in system (6.8). As $\epsilon\to 0$, the path of system (6.7) satisfies the large deviation principle, and we can quantify the probability distribution of the system’s deviation from the most probable path using the rate function.
@@
Figure 5. The comparison figure of Fig 4 when $\epsilon=0.01$ and $\epsilon=0.1$, respectively.
@@
acknowledgments
The reviewers’ suggestions have helped us to perfect the theoretical framework, supplement the high - dimensional numerical experiments, and enhance the originality and depth of the paper. The editor’s support has given us the opportunity to improve the paper. We hereby express our sincere gratitude.

### v2

@@ 6 Preservation of Invariant Tori in Nearly Integrable Stochastic Hamiltonian Systems
$\sup_{D_{r_{j},s_{j}}(\mathscr{D}_{j})}|\text{Im}\phi^{j}|\leq\frac{\xi_{j}}{2}.$ (44)
Hence
@@
Thus, thanks to (43), $\mathcal{H}^{\epsilon}_{j}\circ\phi^{j}=\mathcal{H}^{j}+\mathcal{P}^{j}$ satisfies the assumptions in (3) with $\varepsilon\sim\left\lVert\mathcal{P}^{j}\right\rVert_{r_{j},s_{j},\mathscr{D}_{j}}$, $r\sim r_{j}$, $s\sim s_{j}$, $\sigma\sim\sigma_{j}$, $C_{H}\sim 2C_{H}$ as
@@
Hence, in order to apply Lemma 6 to $\mathcal{H}^{\epsilon}_{j}\circ\phi^{j}=\mathcal{H}^{j}+\mathcal{P}^{j}$, we only need to check (5). Upon observation, we have
@@
Furthermore, based on Lemmas 2.17 and 2.18, we can obtain the convergence results for $G^{j}$, $P_{j}$, $\phi^{j}$, and $H_{j}$ as follows:
  • The sequence $G^{j}:=G_{j}\circ G_{j-1}\circ\cdots\circ G_{2}\circ G_{1}$ converges uniformly on $\mathscr{D}_{0}$ to a diffeomorphism $G_{*}:\mathscr{D}_{0}\to\mathscr{D}_{*}:=G_{*}(\mathscr{D}_{0})\subset\mathscr{D}$, and $G_{*}\in C_{W}^{1}(\mathscr{D}_{0})$;
  • $P_{j}$ converges uniformly to 0 on $\mathscr{D}_{*}\times\mathbb{T}_{s_{*}}^{d}$ in the $C_{W}^{2}$ topology;
@@
    with $\phi_{*}\in C_{W}^{\tilde{m}}(\mathscr{D}_{*}\times\mathbb{T}^{d})$ and $\phi_{*}(\cdot,\cdot)\in C^{\tilde{m}v}(\mathbb{T}^{d})$, for any given $y_{*}\in\mathscr{D}_{*}$;
@@
Based on the preceding proof, we can demonstrate that there exist a Cantor-like set $\mathcal{D}_{*}\subset\mathcal{D}$, an embedding $\phi_{*}=(v_{*},u_{*}):\mathcal{D}_{*}\times\mathbb{T}^{d}\rightarrow\mathcal{H}:=\phi_{*}(\mathcal{D}_{*}\times\mathbb{T}^{d})\subset\mathcal{D}\times\mathbb{T}^{d}$ of class $C_{W}^{\beta}(\mathcal{D}_{*}\times\mathbb{T}^{d})$, and a function $H_{*}\in C_{W}^{2}(\mathcal{D}_{*},\mathbb{R})$, such that $H^{\epsilon}\circ\phi_{*}(I_{*},\theta)=K_{*}(I_{*},\theta)$ for all $(I_{*},\theta)\in\mathcal{D}_{*}\times\mathbb{T}^{d}$. The map $\xi\mapsto\phi_{*}(I_{*},\theta)$ is of class $C_{\beta}^{v}(\mathbb{T}^{d})$ for any $I_{*}\in\mathcal{D}_{*}$ (with $v^{-1}<\beta<\beta_{0}$), and the map $G^{*}:=(\partial_{I_{*}}H_{*})^{-1}\circ\partial_{I}H:\mathcal{D}_{*}\rightarrow\mathcal{D}_{*}$ defines a lipeomorphism onto $\mathcal{D}_{*}$, satisfying $B_{\alpha/2}(\mathcal{D}_{*})\subseteq\mathcal{D}$. The set $\mathcal{K}$ is foliated by KAM tori of $H^{\epsilon}$, each being a graph of a $C^{v}(\mathbb{T}^{d})$-map.
@@
$\epsilon^{2}\ln\mathbb{P}(X(t)\in A)\approx-\inf_{\varphi\in A}S(\varphi),$
where $\varphi-x_{0}\in\mathbb{H}^{1}$, $A\in\mathbb{R}^{d}$ denotes an arbitrary measurable set and the rate function $S(\varphi)$ is given by
$S(\varphi)$ $=\frac{1}{2}\left(\int_{0}^{1}\left\|\sigma^{-1}(s,\varphi_{s})\left(J\nabla H_{\epsilon}(\varphi_{s})-\dot{\varphi_{s}}\right)\right\|^{2}\,{\rm d}s\right).$
@@
$\mathbb{P}(X_{\epsilon}(t)\in\mathcal{K}_{0})\approx\exp\left\{-C\frac{\epsilon_{1}^{2}}{\epsilon_{2}^{2}}\right\}\times\mathbb{P}(\left\|W^{\sigma}\right\|\leq 1).$
@@
Based on the content of Theorem 6.1, we have
$\quad\frac{\mathbb{P}(X_{\epsilon}(t)\in\mathcal{K}_{0})}{\mathbb{P}(\left\|W^{\sigma}\right\|\leq 1}\approx\exp\left\{-\frac{1}{\epsilon_{2}^{2}}\inf_{\varphi\in\mathcal{K}_{0}}S(\varphi)\right\}$
$=\exp\left\{-\frac{1}{2\epsilon_{2}^{2}}\inf_{\varphi\in\mathcal{K}_{0}}\left(\int_{0}^{1}\left\|\sigma^{-1}(t)\left(\epsilon_{1}J\nabla P(\varphi_{s})\right)\right\|^{2}\,{\rm d}s\right)\right\}$
$=\exp\left\{-\frac{\epsilon_{1}^{2}}{2\epsilon_{2}^{2}}\inf_{\varphi\in\mathcal{K}_{0}}\left(\int_{0}^{1}\left\|\sigma^{-1}(t)\left(\epsilon_{1}J\nabla P(\varphi_{s})\right)\right\|^{2}\,{\rm d}s\right)\right\}$
@@
where $C$ is a quantity that depends on $\sigma^{-1}$ and the partial derivatives of $P$. ∎
@@
[…]
@@
[…]
@@
The Hamiltonian of this system can be written as
@@
From the theorems discussed earlier, we know that the most probable continuous path of the stochastic nearly integrable Hamiltonian system (45) is governed by the deterministic equations in system (46). As $\epsilon\to 0$, the path of system (45) satisfies the large deviation principle, and we can quantify the probability distribution of the system’s deviation from the most probable path using the rate function.
@@
Figure 3: The comparison figure of Fig 2 when $\epsilon=0.01$ and $\epsilon=0.1$, respectively.